A garden measuring 70 feet by 60 feet is to have a pedestrian walkway built around all four of its sides. The total area of the garden, including the walkway, will be 5690.36 ft^2. What is the width of the walkway?
Answer by richwmiller(17219)

(70+x)*(60+x)=5960.36
x^2+130 x+4200 = 5960.36
x^2+130 x-1760.36 = 0
Usable answer
x = 12.3651